# Categorical Syllogisms Revisited: A Review of the Logical Reasoning Abilities of LLMs for Analyzing Categorical Syllogisms

### v1

@@ Abstract
There have been a huge number of benchmarks proposed to evaluate how large language models (LLMs) behave for logic inference tasks. However, it remains an open question how to properly evaluate this ability. In this paper, we provide a systematic overview of prior works on the logical reasoning ability of LLMs for analyzing categorical syllogisms. We first investigate all the possible variations for the categorical syllogisms from a purely logical perspective and then examine the underlying configurations (i.e., mood and figure) tested by the existing datasets. Our results indicate that compared to template-based synthetic datasets, crowdsourcing approaches normally sacrifice the coverage of configurations (i.e., mood and figure) of categorical syllogisms for more language variations, thus bringing challenges to fully testing LLMs under different situations. We then proceed to summarize the findings and observations for the performances of LLMs to infer the validity of syllogisms from the current literature. The error rate breakdown analyses suggest that the interpretation of the quantifiers seems to be the current bottleneck that limits the performances of the LLMs and is thus worth more attention. Finally, we discuss several points that might be worth considering when researchers plan on the future release of categorical syllogism datasets. We hope our work will not only provide a timely review of the current literature regarding categorical syllogisms, but also motivate more interdisciplinary research between communities, specifically computational linguists and logicians. [^1] Preprint. Under review.
@@
Shi Zong, Jimmy Lin University of Waterloo {s4zong, jimmylin}@uwaterloo.ca
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performances on a variety of tasks Brown et al. (2020); Wei et al. (2022); Bubeck et al. (2023). Over the years, a large number of benchmarks have been proposed that try to evaluate the different abilities of LLMs, many of which are designed for measuring logical reasoning ability using a variety of tasks. Habernal et al. (2018) propose an argument reading comprehension task to test deductive reasoning. CLUTRR Sinha et al. (2019) tests inductive reasoning capabilities by requiring to infer kinship relations between characters in short stories. ReClor Yu et al. (2020), MMLU Hendrycks et al. (2021), and LogiQA Liu et al. (2020) contain multiple-choice reading comprehension questions to evaluate diverse forms of logical reasoning. Datasets such as SylloBase Wu et al. (2023) and FOLIO Han et al. (2022) require LLMs to conduct inferences using syllogism logic or first-order logic. Among these datasets, many consist of questions that are directly taken from exams. For example, MMLU Hendrycks et al. (2021) contains practice questions from tests such as the Graduate Record Examination (GRE), and ReClor Yu et al. (2020) collects problems from the Law School Admission Test (LSAT).
@@
A fundamental question behind these datasets is: how to design a benchmark to ensure a fair and comprehensive evaluation of logic reasoning abilities. This question is particularly important when the test questions are self-generated, instead of directly collected from established examinations for humans mentioned above. Problems in those human examinations are developed over decades and are designed in support of theories such as psychometrics and measurement in education. Thus, having rigorous analyses of current benchmarks designed for LLMs would ensure that we can track the development progress of LLMs accurately.
@@
In this work, we make progress in answering the above question for a specific task: categorical syllogisms. [^2] Unless specified, the term “categorical syllogisms” is also directly written as “syllogisms” (due to space issues). Besides the reason that to the best of our knowledge, there is no prior work on analyzing categorical syllogism datasets from a designing principle’s perspective, we note some other compelling reasons for choosing this task. (1) Syllogism is inarguably the most basic building block in logical reasoning abilities. Having a deeper understanding of syllogism inference is thus beneficial when designing models for solving more complex reasoning tasks. (2) Categorical syllogisms have a finite number of situations (discussed in Section 2.1), which could enable a complete check of all the possible cases for LLMs. (3) How to properly solve categorical syllogisms has been studied by logicians over decades. There is a huge literature that we can draw inspiration from to help understand how LLMs behave or make LLMs more efficient.
@@
To sum up, our intention is not to propose new models to achieve the start-of-the-performances on certain datasets, nor introduce new benchmarks. Rather, we hope to take a step back and systematically review all existing work to understand where we are right now. Our goal is to check missing pieces and identify areas that are worth clarifying or need future research. Specifically, in this paper, we make the following contributions:
@@
  • We investigate all the existing categorical syllogism datasets in literature along with their properties in Section 3. A checklist covering all the variations of categorical syllogisms from a purely logician’s perspective is provided and we then examine the coverage of different cases for existing benchmarks.
@@
  • We summarize all prior findings related to the performances of LLMs for checking the validity of syllogisms in Section 4. By presenting an error rate breakdown by the mood and figure of syllogisms, we highlight the importance of enhancing the abilities of LLMs for interpreting quantifiers.
@@ 2 A Concise Introduction to Syllogisms
In this section, we provide a brief introduction to the categorical syllogisms from a logician’s perspective. We will show in Sections 3.2 and 4.3 that these preparations will help us evaluate the current syllogism datasets and better understand the bottleneck of the performances of LLMs.
@@ 2.1 Categorical Syllogisms
Since there are 4 kinds of categorical propositions and 3 categorical propositions in a categorical syllogism, there are 64 possible moods ($644^{3}=$ 64). As each mood can occur in each of the four figures, in total we have $2564^{4}=$ 256 different syllogisms. Among these, only 24 are valid forms, which are extensively studied by the logicians.33315 configurations are “unconditionally valid” and another 9 are “conditionally valid”. It is related to existential import in Section 5.1. Thus, we have the following fact: the validity of the standard syllogism can be determined by checking the configuration (mood and figure) against a list of valid syllogistic forms.
@@ 2.2 Analyzing Syllogisms as a Logician
In practice, rare propositions are in their standard form and we need to make translations. The major benefit of such translation is that the operations and inferences pertinent to standard-form categorical propositions can be directly applied to these statements. Logicians have developed a number of well-tested methods for translating nonstandard propositions, although given the richness of ordinary language, these specific rules can not cover all possible cases.
@@
Logical Human authored questions     valid Conclusion valid- PaLM 2-L $\sim$9048 No
Lampinen et al. (2023) belief-consistent -ity identification (support)
NeuBAROCO BAROCO (originally designed     entail, contra, neu Conclusion valid- GPT-3.5 51.7% 375 No
Ando et al. (2023) for human intell. test) inference types -ity identification (overall)
@@
FOLIO Template w/ cro- N/A true, false, Conclusion truth Logic-LM 78.1% 1,435 Yes
Han et al. (2022) wdsourcing rewrite unknown identification (GPT-4)
@@
Table 4: Overview of existing syllogism datasets, along with their construction methods, annotations included, and the documented model performances. We also list two representative datasets that are designed for testing first-order logic but involve syllogisms (discussed in Section 3).
@@
3 Reviewing Existing Syllogism Datasets
@@ 3.1 Summary of Syllogism Datasets
Templated based Approach.
@@
Text Generated by Human.
Non-synthetic datasets are normally developed through crowdsourcing efforts. To acquire high-quality inference questions efficiently, these datasets sometimes rely on guidance during the crowdsourcing tasks. SylloFigure Peng et al. (2021) is built based on the idea of enthymeme reconstruction. Specifically, Peng et al. (2021) select the entailment part of SNLI Bowman et al. (2015) dataset and then add the annotations of figures. Avicenna Aghahadi and Talebpour (2022) is a crowdsourcing dataset, and the syllogisms are extracted from various sources, such as books and news articles. Syllogisms in Lampinen et al. (2023) are hand-authored. NeuBAROCO Ando et al. (2023) originates from BAROCO, which is written in Japanese and is developed to evaluate human syllogistic reasoning abilities. FOLIO Han et al. (2022) first generates logically valid stories using syllogism templates and then asks human annotators to write logically-valid stories in natural language.
@@
As shown in Table 4, nearly all datasets with human-generated text lack certain kinds of annotations, thus causing troubles to analyze them (in Section 3.3). We fill in this missing gap by collecting relevant examples and corresponding exercise questions from standard introduction to logic textbooks Copi et al. (2019); Kelley (2013); Baronett (2018); Hurley and Watson (2018).
In total, we collect 371 examples of translating statements into standard-form, covering all the possible forms of phraseology discussed in Section 3.2; 64 examples for judging the types of standard propositions; and 116 examples for judging the validity of a given syllogism, with complete annotations for the mood and figure. Among these examples, 57 are enthymemes. All these examples will be released publicly.
@@ 3.2 Variations of Categorical Syllogisms
A set of questions that cover all the possible cases could be achieved by varying components in different levels of granularity that we outline in Section 2.1. We consider all possible variations from two angles: syllogisms in standard and non-standard forms. For standard syllogisms, the underlying nature is decided by the combination of mood and figure, which leads to 256 different cases.
For non-standard syllogism, there are variations both on the individual proposition level and the syllogism level. On the proposition level, we consider the different options of quantifiers, terms, and copula: (1) Besides standard quantifiers, the propositions could have non-standard quantifiers (also known as generalized quantifiers), such as “few”, “a few”, “not every”, or “anyone”, and unexpressed quantifiers; (2) Terms could be expressed with only an abjective, a plural noun or a pronoun, and the verbs are in other forms of the verb “to be;” and (3) Certain propositions could be typically translated into categorical propositions. Established categories include singular propositions, conditional statements such as “if … then,” exclusive propositions that involve words “only,” “none but,” and “none except,” and exceptive propositions in the form of “All except S are P” and “All but S are P”.
@@ 3.3.1 Building Tools for Assessing Coverage
Most of the datasets do not have the annotations needed to be assessed (details in Table 4). Making up these missing pieces would require human annotators with linguistic background. Given the huge amount of human effort for such annotations, we take the approach of directly asking LLMs to perform as an annotator for labeling.
To ensure that we can build prompts with reasonable performances, we calibrate them on our newly collected textbook questions (discussed in Section 3.1). We also use the fact about the validity of syllogisms mentioned in Section 2.1 for cross-checking: for a valid inference, if a predicted configuration is not one of the valid syllogism forms, then there is something wrong with this prediction.
@@
When translating statements, besides a deep understanding of the given statement, we need to follow some established rules set by the logicians (for example, the treatment of singular propositions discussed in Appendix A). We thus base our prompt design on a 2-step translation process: (1) determine the nature of a proposition by classifying it into categories list in Table 5), and (2) then perform the translation based on the set rules within that category. To make sure the translated proposition is in the standard form, we also set up a mechanism for a second-round translation. We observe that GPT-4o performs well in identifying the forms of phraseology, while it is easy to incorrectly classify some statements into singular propositions. A manual check for the translated propositions shows that GPT-4 achieves 87.3% accuracy on 371 textbook problems, with 68 propositions translated twice.
@@
We may not first translate individual propositions and then simply combine the detected proposition types together to form the mood of the syllogism, due to the issue of having potentially more than three terms (in Appendix A). Thus, we feed the syllogism as a whole and ask GPT-4 to generate the mood and figure simultaneously. The principles and rules discussed above for translating propositions are also incorporated into the prompt. Experimental results on 116 textbook examples reveal an accuracy of 87.9% for mood detection, 48.3% for figure detection, and 44.8% for configuration detection. A further review of mood detection results reveals that this high accuracy is due to the fact that most of the collected textbook examples are standard-form propositions.
@@ 3.3.2 Datasets Coverage Observations
Our assessment results are reported in Table 5 (detailed configurations are in Figure 1). In Figure 2 we also provide the distribution of the estimated proposition types in the SylloFigure dataset. We observe that both the proposition types (A, E, I, O) and forms of phraseology (types in Table 5) are distributed highly unevenly, and datasets normally have different distributions. Regarding the coverage of configurations, we observe that compared to template-based datasets, datasets using human-generated text are normally centered on few specific moods and figures, i.e., Avicenna covers only over 7 different syllogisms configurations, calculated from 60.9% of the whole dataset.
Since we use LLMs instead of human efforts to make up the missing mood and figures, the coverage percentages in Table 5 could only be treated as rough estimates. Nevertheless, our key point is clear: datasets that are from crowdsourcing efforts are skewed to certain linguistic styles and cover only limited configurations of syllogisms. We thus suggest researchers take the actual variations covered by the datasets into account when interpreting experimental results.
@@ 4.1 What Do We Know So Far?
We observe prior studies mainly make use of the following approaches to evaluate the validity of categorical syllogisms: (1) given two premises, select a correct conclusion from multiple choices Wu et al. (2023); Eisape et al. (2024), (2) given two premises and a conclusion, identify if the logic inference is valid Lampinen et al. (2023) or entail Ando et al. (2023), and (3) given two premises or more, generate the conclusion Aghahadi and Talebpour (2022); Saparov and He (2023). In general, most prior works report LLMs have an accuracy of around 75% when evaluating the validity of given syllogisms. We provide more performance evaluation details in Table 4.
@@
4.2 LLMs’ Performances Breakdowns by Syllogisms Configurations

### v2

@@ Abstract
There has been a huge number of benchmarks proposed to evaluate how large language models (LLMs) behave for logic inference tasks. However, it remains an open question how to properly evaluate this ability. In this paper, we provide a systematic overview of prior works on the logical reasoning ability of LLMs for analyzing categorical syllogisms. We first investigate all the possible variations for categorical syllogisms from a purely logical perspective and then examine the underlying configurations (i.e., mood and figure) tested by existing datasets. Our results indicate that compared to template-based synthetic datasets, crowdsourcing approaches normally sacrifice the coverage of configurations (i.e., mood and figure) of categorical syllogisms for more language variations, thus bringing challenges to fully testing LLMs under different situations. We then summarize the findings and observations for the performance of LLMs to infer the validity of syllogisms from the current literature. The error rate breakdown analyses suggest that the interpretation of quantifiers seems to be the current bottleneck that limits the performance of the LLMs and is thus worth more attention. Finally, we discuss several points that might be worth considering when researchers plan to release categorical syllogism datasets. We hope our work will provide a timely review of the current literature regarding categorical syllogisms, and motivate more interdisciplinary research between communities, specifically computational linguists and logicians.
@@
Shi Zong, Jimmy Lin David R. Cheriton School of Computer Science University of Waterloo {s4zong, jimmylin}@uwaterloo.ca
{CJK*}
UTF8gbsn
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance on a variety of tasks Brown et al. (2020); Wei et al. (2022); Bubeck et al. (2023). Over the years, a large number of benchmarks have been proposed that try to evaluate the different abilities of LLMs, many of which are designed for measuring logical reasoning ability using a variety of tasks. Habernal et al. (2018) propose an argument reading comprehension task to test deductive reasoning. CLUTRR Sinha et al. (2019) tests inductive reasoning capabilities by requiring to infer kinship relations between characters in short stories. ReClor Yu et al. (2020), MMLU Hendrycks et al. (2021), and LogiQA Liu et al. (2020) contain multiple-choice reading comprehension questions to evaluate diverse forms of logical reasoning. Datasets such as SylloBase Wu et al. (2023) and FOLIO Han et al. (2022) require LLMs to conduct inferences using syllogism logic or first-order logic. Among these datasets, many consist of questions that are directly taken from exams. For example, MMLU Hendrycks et al. (2021) contains practice questions from tests such as the Graduate Record Examination (GRE), and ReClor Yu et al. (2020) collects problems from the Law School Admission Test (LSAT).
@@
A fundamental question behind these datasets is: how to design a benchmark to ensure a fair and comprehensive evaluation of logic reasoning abilities? This question is particularly important when the test questions are self-generated, instead of directly collected from established examinations for humans mentioned above. Problems in those human examinations are developed over decades and are designed in support of theories such as psychometrics and measurement in education. Thus, having rigorous analyses of current benchmarks designed for LLMs would ensure that we can track the development progress of LLMs accurately.
@@
In this work, we make progress in answering the above question for a specific task: categorical syllogisms. [^1] Unless specified, the term “categorical syllogisms” is also directly written as “syllogisms” (due to space issues). Besides the reason that to the best of our knowledge, there is no prior work on analyzing categorical syllogism datasets from a designing principle’s perspective, we note some other compelling reasons for choosing this task. (1) Syllogisms are inarguably the most basic building block in logical reasoning abilities. Having a deeper understanding of syllogism inference is thus beneficial when designing models for solving more complex reasoning tasks. (2) Categorical syllogisms have a finite number of situations (discussed in Section 2.1), which could enable a complete check of all the possible cases for LLMs. (3) How to properly solve categorical syllogisms has been studied by logicians over decades. There is a huge literature that we can draw inspiration from to help understand how LLMs behave or make LLMs more efficient.
@@
To sum up, our intention is not to propose new models to achieve the start-of-the-performance on certain datasets, nor introduce new benchmarks. Rather, we hope to take a step back and systematically review all existing work to understand where we are right now. Our goal is to check missing pieces and identify areas that are worth clarifying or need future research. Specifically, in this paper, we make the following contributions:
@@
  • We investigate all existing categorical syllogism datasets in literature along with their properties in Section 3. A checklist covering all the variations of categorical syllogisms from a purely logician’s perspective is provided and we then examine the coverage of different cases for existing benchmarks.
@@
  • We summarize all prior findings related to the performance of LLMs for checking the validity of syllogisms in Section 4. By presenting an error rate breakdown by the mood and figure of syllogisms, we highlight the importance of enhancing the abilities of LLMs for interpreting quantifiers.
@@ 2 A Concise Introduction to Syllogisms
In this section, we provide a brief introduction to categorical syllogisms from a logician’s perspective. We will show in Sections 3.2 and 4.3 that these preparations will help us evaluate current syllogism datasets and better understand the bottleneck of the performance of LLMs.
@@ 2.1 Categorical Syllogisms
Since there are 4 kinds of categorical propositions and 3 categorical propositions in a categorical syllogism, there are 64 possible moods ($644^{3}=$ 64). As each mood can occur in each of the four figures, in total we have $2564^{4}=$ 256 different syllogisms. Among these, only 24 are valid forms, which are extensively studied by logicians.22215 configurations are “unconditionally valid” and another 9 are “conditionally valid”. It is related to existential import in Section 5.1. Thus, we have the following fact: the validity of the standard syllogism can be determined by checking the configuration (mood and figure) against a list of valid syllogistic forms.
@@ 2.2 Analyzing Syllogisms as a Logician
In practice, rare propositions are in their standard form and we need to make translations. The major benefit of such translation is that the operations and inferences pertinent to standard-form categorical propositions can be directly applied to these statements. Logicians have developed a number of well-tested methods for translating non-standard propositions, although given the richness of ordinary language, these specific rules can not cover all possible cases.
@@
Logical Human authored questions     valid Conclusion validity PaLM 2-L $\sim$9048 No
Lampinen et al. (2023) belief-consistent identification (support)
NeuBAROCO BAROCO (originally designed     entail, contra, neu Conclusion validity GPT-3.5 51.7% 375 No
Ando et al. (2023) for human intell. test) inference types identification (overall)
@@
FOLIO Template w/ crowd N/A true, false, Conclusion truth Logic-LM 78.1% 1,435 Yes
Han et al. (2022) -sourcing rewrite unknown identification (GPT-4)
@@
Table 4: Overview of existing syllogism datasets, along with their construction methods, annotations included, and the documented model performance.
@@
3 Review of Existing Syllogism Datasets
@@ 3.1 Summary of Syllogism Datasets
Template-based Approach.
@@
Text Generated by Humans.
Non-synthetic datasets are normally developed through crowdsourcing efforts. To acquire high-quality inference questions efficiently, these datasets sometimes rely on guidance during the crowdsourcing tasks. SylloFigure Peng et al. (2021) is built based on the idea of enthymeme reconstruction. Specifically, Peng et al. (2021) select the entailment part of the SNLI Bowman et al. (2015) dataset and then add the annotations of figures. Avicenna Aghahadi and Talebpour (2022) is a crowdsourcing dataset, and the syllogisms are extracted from various sources, such as books and news articles. Syllogisms in Lampinen et al. (2023) are hand-authored. NeuBAROCO Ando et al. (2023) originates from BAROCO, which is written in Japanese and is developed to evaluate human syllogistic reasoning abilities. FOLIO Han et al. (2022) first generates logically valid stories using syllogism templates and then asks human annotators to write logically valid stories in natural language.
@@
As shown in Table 4, nearly all datasets with human-generated text lack certain kinds of annotations, thus causing troubles in analyzing them (in Section 3.3). We fill in this missing gap by collecting relevant examples and corresponding exercise questions from standard introduction to logic textbooks Copi et al. (2019); Kelley (2013); Baronett (2018); Hurley and Watson (2018).
In total, we collect 371 examples of translating statements into standard form, covering all the possible forms of phraseology discussed in Section 3.2; 64 examples for judging the types of standard propositions; and 116 examples for judging the validity of a given syllogism, with complete annotations for the mood and figure. Among these examples, 57 are enthymemes.
@@ 3.2 Variations of Categorical Syllogisms
A set of questions that cover all the possible cases could be achieved by varying components of different levels of granularity that we outline in Section 2.1. We consider all possible variations from two angles: syllogisms in standard and non-standard forms. For standard syllogisms, the underlying nature is decided by the combination of mood and figure, which leads to 256 different cases.
For non-standard syllogism, there are variations both on the individual proposition level and the syllogism level. On the proposition level, we consider the different options of quantifiers, terms, and copula: (1) Besides standard quantifiers, the propositions could have non-standard quantifiers (also known as generalized quantifiers), such as “few”, “a few”, “not every”, or “anyone”, and unexpressed quantifiers; (2) Terms could be expressed with only an adjective, a plural noun or a pronoun, and the verbs are in other forms of the verb “to be;” and (3) Certain propositions could be typically translated into categorical propositions. Established categories include singular propositions, conditional statements such as “if … then,” exclusive propositions that involve words “only,” “none but,” and “none except,” and exceptive propositions in the form of “All except S are P” and “All but S are P”.
@@ 3.3.1 Building Tools for Assessing Coverage
Most of the datasets do not have the annotations needing to be assessed (details in Table 4). Making up these missing pieces would require human annotators with linguistic background. Given the huge amount of human effort for such annotations, we take the approach of directly asking LLMs to perform as an annotator for labeling.
To ensure that we can build prompts with reasonable performance, we calibrate them on our newly collected textbook questions (discussed in Section 3.1). We also use the fact about the validity of syllogisms mentioned in Section 2.1 for cross-checking: for a valid inference, if a predicted configuration is not one of the valid syllogism forms, then there is something wrong with this prediction.
@@
When translating statements, besides a deep understanding of the given statement, we need to follow some established rules set by logicians (for example, the treatment of singular propositions discussed in Appendix A). We thus base our prompt design on a 2-step translation process: (1) determine the nature of a proposition by classifying it into categories listed in Table 5), and (2) then perform the translation based on the set rules within that category. To make sure the translated proposition is in the standard form, we also set up a mechanism for a second-round translation. We observe that GPT-4o performs well in identifying the forms of phraseology, while it is easy to incorrectly classify some statements into singular propositions. A manual check for the translated propositions shows that GPT-4 achieves 87.3% accuracy on 371 textbook problems, with 68 propositions translated twice.
@@
We can not first translate individual propositions and then simply combine the detected proposition types together to form the mood of the syllogism, due to the issue of having potentially more than three terms (in Appendix A). Thus, we feed the syllogism as a whole and ask GPT-4 to generate the mood and figure simultaneously. The principles and rules discussed above for translating propositions are also incorporated into the prompt. Experimental results on 116 textbook examples reveal an accuracy of 87.9% for mood detection, 48.3% for figure detection, and 44.8% for configuration detection. A further review of mood detection results reveals that this high accuracy is due to the fact that most of the collected textbook examples are standard-form propositions.
@@ 3.3.2 Datasets Coverage Observations
Our assessment results are reported in Table 5 (detailed configurations are in Figure 1). In Figure 2 we also provide the distribution of the estimated proposition types in the SylloFigure dataset. We observe that both the proposition types (A, E, I, O) and forms of phraseology (types in Table 5) are distributed highly unevenly, and datasets normally have different distributions. Regarding the coverage of configurations, we observe that compared to template-based datasets, datasets using human-generated text are normally centered on a few specific moods and figures, i.e., Avicenna covers only over 7 different syllogisms configurations, calculated from 60.9% of the whole dataset.
Since we use LLMs instead of human effort to make up the missing mood and figures, the coverage percentages in Table 5 can only be treated as rough estimates. Nevertheless, our key point is clear: datasets that are from crowdsourcing efforts are skewed to certain linguistic styles and cover only limited configurations of syllogisms. We thus suggest researchers take the actual variations covered by the datasets into account when interpreting experimental results.
@@ 4.1 What Do We Know So Far?
We observe prior studies mainly make use of the following approaches to evaluate the validity of categorical syllogisms: (1) given two premises, select a correct conclusion from multiple choices Wu et al. (2023); Eisape et al. (2024), (2) given two premises and a conclusion, identify if the logic inference is valid Lampinen et al. (2023); Ando et al. (2023), and (3) given two premises or more, generate the conclusion Aghahadi and Talebpour (2022); Saparov and He (2023). In general, most prior works report LLMs have an accuracy of around 75% when evaluating the validity of given syllogisms. We provide more performance evaluation details in Table 4.
@@
4.2 LLMs’ Performance Breakdowns by Syllogisms Configurations